# Vertex F-algebras and their associated Lie algebra

### v1

@@
Abstract
Vertex $F$-algebras are a deformation of the concept of an ordinary vertex algebra in which the additive formal group law is replaced by an arbitrary formal group law $F$. The main theorem of this paper constructs a Lie algebra from a vertex $F$-algebra — for the additive formal group law, this extends Borcherds’ well-known construction for ordinary vertex algebras. Our construction involves the new concept of an $F$-residue and some other new algebraic concepts, which are deformations of familiar concepts for the special case of an additive formal group law.
@@ 1 Introduction and Results
A fundamental theorem of Borcherds [1], which has important applications to representation theory, states that every vertex algebra determines a Lie algebra. The main result of this paper extends this result and shows that every vertex $F$-algebra also determines a Lie algebra. However, the usual construction of the Lie bracket in the vertex algebra setting, via the residue of the state-to-field correspondence, does not directly apply to vertex $F$-algebras. To overcome this difficulty, we introduce a deformation of the concept of a residue, the $F$-residue $\operatorname{Res}_{z=0}^{F}$. Establishing the Lie algebra structure in the generalized setting requires significantly more technical work than in the classical case. The following result is our main theorem, and it is proven in §4.3.
@@
Let $(V,\mathbb{1},\mathcal{D},Y)$ be a vertex $F$-algebra (see Definition 4.1). Then
$[a,b]=\operatorname{Res}_{z=0}^{F}Y(a,z)b\,dz.$
defines a Lie bracket on the quotient $V/\sum_{n\geqslant 1}\mathcal{D}^{(n)}(V).$
Besides the potential applications to representation theory, there is further motivation for generalizing the concept of a vertex algebra stemming from Enumerative Geometry: recently, Joyce [8, 9] has shown that vertex algebras play a central role in Enumerative Geometry, where the Lie algebra associated to the ordinary homology $H_{*}(\mathcal{M})$ of a moduli stack $\mathcal{M}$ (for example, the stack of perfect complexes over a projective variety). The Lie bracket is used to formulate wall-crossing formulas of enumerative invariants. The study of enumerative invariants in generalized homology $E_{*}(\mathcal{M})$ is becoming increasingly popular (most notably, the theory of $K$-theoretic enumerative invariants). For complex oriented generalized homology theories $E_{*}$ with formal group law $F$, similar constructions show that $E_{*}(\mathcal{M})$ is naturally a vertex $F$-algebra. Indeed, in [5] the author and J. Gross proved that the generalized homology of an $H$-space (plus some extra data) naturally has a vertex $F$-algebra structure. This extends Joyce’s construction to generalized homology. While a comprehensive theory of wall-crossing has not yet been developed for generalized enumerative invariants (for progress in the case of $K$-theory, see Liu [14]), it may well be expected that the wall-crossing formula will use the Lie bracket constructed in this paper.
The proof of our main theorem, as well as a better understanding of vertex $F$-algebras, requires the development of various deformations of familiar concepts ($F$-binomial coefficients, $F$-delta distributions, $F$-residues, and $F$-hyperderivatives), developed in §3. These constructions, which appear to be new to the literature, provide the tools for generalizing results for ordinary vertex algebras to vertex $F$-algebras. §2 reviews some background of formal groups and establishes certain properties we will use later. In §4 we define vertex $F$-algebras (our definition is equivalent to that of Li [13]), prove some meromorphicity properties, and prove our main theorem.
Notation.
[…]
  • $R[z^{\pm 1}]$ the ring of Laurent polynomials
  • $R\llbracket z\rrbracket$ the ring of formal power series $\sum_{i=0}^{\infty}a_{i}z^{i}$
  • $R\llbracket z^{\pm 1}\rrbracket$ the group of bilateral Laurent series $\sum_{i=-\infty}^{+\infty}a_{i}z^{i}$ (note that $R\llbracket z^{\pm 1}\rrbracket$ is not a ring, but only has a partially defined product)
  • $R\llparenthesis z\rrparenthesis$ the ring of meromorphic series having a pole at 00
@@ 2 Background on Formal Groups
$F(z,w)$ $=F(w,z)=z+w+O(zw),$ $F(z,F(w,v))$ $=F(F(z,w),v).$ (2.1)
There is a unique inverse $\iota(z)=-z+O(z^{2})\in R\llbracket z\rrbracket$ such that $F(z,\iota z)=0.$
@@
Over $R=\mathbb{Z}[s]$ we have the formal group law
@@
We view $F_{s}$ as a family of group laws over the affine line. At $s=0$ it specializes to the additive group law $F_{a}$ and at $s=1$ to the multiplicative group law $F_{m}.$ All polynomial formal group laws are of the form (2.2), and other examples must involve infinite series.
@@
Over the ring of modular forms $R=\mathbb{Z}[\frac{1}{2},\epsilon,\delta]$ we have the elliptic formal group law
$F_{ell}(z,w)=\frac{z\sqrt{S(w)}+w\sqrt{S(z)}}{1-\epsilon z^{2}w^{2}},\quad S(z)=1-2\delta z^{2}+\epsilon z^{4}.$
For $\epsilon=\delta^{2}$ it specializes to the group law
$\frac{z+w}{1+\delta zw}$
(for $\delta=1$ equivalent to $L$-genus) and for $\epsilon=0$ it specializes to
@@
Let $p$ be a prime number and fix a positive power $q=p^{h}.$ Set $\phi_{p}(z)=z+\sum_{n\geqslant 1}p^{-n}X^{q^{n}}.$ In [6] it is shown that
@@
$\theta_{F}=p_{F}(z)dz,\qquad p_{F}(z)\in R\llbracket z\rrbracket,$ (2.4)
satisfying
$p_{F}(0)=1,\qquad F^{*}(\theta_{F})=\pi_{1}^{*}(\theta_{F})+\pi_{2}^{*}(\theta_{F}).$ (2.5)
Equivalently,
$p_{F}(F(z,w))\bigl{(}F^{1,0}(z,w)dz+F^{0,1}(z,w)dw\bigr{)}=p_{F}(z)dz+p_{F}(w)dw.$ (2.6)
@@
$-\iota^{*}(\theta_{F})=\theta_{F}.$ (2.7)
@@
$\phi(F(z,w))$ $=\phi(z)+\phi(w),$ $\phi(0)$ $=0,$ $\phi^{\prime}(0)$ $=1.$ (2.8)
Indeed, taking $\partial/\partial w$ of (2.8) and setting $w=0$ we find that $\phi^{\prime}(z)=\theta_{F}$ is the invariant $1$-form which, since $\mathbb{Q}\subset R,$ has a primitive $\theta_{F}=d\phi.$ The composition inverse $\phi^{-1}(x)\in R\llbracket x\rrbracket$ is called the exponential.
@@
[…]
@@
The logarithm in Example 2.4 is the series $\phi(z)$ defined there. For $p=q$ we have
$p_{F}(z)^{-1}=F^{0,1}(z,0)=\frac{1}{\phi^{\prime}(z)}=\prod_{k=1}^{\infty}\frac{1}{1+z^{p^{k}}}$
Hence the coefficient of $z^{d}$ in $p_{F}(z)^{-1}$ is the sum of $(-1)^{k}$ over all partitions $p^{n_{1}}+\ldots+p^{n_{k}}=d$ into positive powers of the prime $p.$
@@
$L\cong\mathbb{Z}[p_{1},p_{2},\cdots]$ is a polynomial ring over the integers. In particular, $L$ is torsion-free.
@@
Substituting $w=z,$ the first factor is $\psi(0)=1$ and the second is $\phi^{\prime}(z).$
3 Formal Calculus and F-Residues
@@ 3.1 Bilateral and Laurent series
Let $R\llbracket z_{1}^{\pm 1},\ldots,z_{n}^{\pm 1}\rrbracket$ be the space of bilateral Laurent series
$f=\sum_{i_{1},\ldots,i_{n}\in\mathbb{Z}}a_{i_{1},\ldots,i_{n}}z_{1}^{i_{1}}\cdots z_{n}^{i_{n}}$ (3.1)
in variables $z_{1},\ldots,z_{n}.$ This is an abelian group under addition, but the product is only partially defined. The product of (3.1) with
$g=\sum_{j_{1},\ldots,j_{n}\in\mathbb{Z}}b_{j_{1},\ldots,j_{n}}z_{1}^{j_{1}}\cdots z_{n}^{j_{n}}$ (3.2)
@@
$fg=\sum_{k_{1},\ldots,k_{n}\in\mathbb{Z}}\Biggl{(}\sum_{i_{1}+j_{1}=k_{1},\cdots,i_{n}+j_{n}=k_{n}}a_{i_{1},\ldots,i_{n}}b_{j_{1},\ldots,j_{n}}\Biggr{)}z_{1}^{k_{1}}\cdots z_{n}^{k_{n}}$ (3.3)
reduces to a sum with only finitely many non-zero terms. By [12, p.24], this product is associative if all products $fg,$ $gh,$ $(fg)h,$ $f(gh)$ converge and the triple product converges.
@@
We say that (3.1) is a formal Laurent series if $a_{i_{1},\ldots,i_{n}}=0$ for all but finitely many negative indices. The subspace of formal Laurent series
$R\llparenthesis z_{1},\ldots,z_{n}\rrparenthesis$
is a ring since all products converge. For $f\in R\llparenthesis z_{1},\ldots,z_{n}\rrparenthesis$ we say also that $f$ is meromorphic in the variables $z_{1},\ldots,z_{n}.$ We have the subring
$R\llbracket\underline{z}_{1},\ldots,\underline{z}_{n}\rrbracket$
of formal power series which are also said to be holomorphic in $\underline{z}_{1},\ldots,\underline{z}_{n}.$ It is useful to underline holomorphic variables in some contexts below.
@@
This generalizes to series with coefficients in an $R$-module $M.$ Then $R\llparenthesis z_{1},\ldots,z_{n}\rrparenthesis$ acts on $M\llparenthesis z_{1},\ldots,z_{n}\rrparenthesis.$ We leave this extension to the reader.
@@
Let $f=\sum_{n\geqslant N}a_{n}z^{n}\in R\llparenthesis z\rrparenthesis$ with lowest coefficient $a_{N}\neq 0.$ Then $f$ is invertible in $R\llparenthesis z\rrparenthesis$ $\iff$ $a_{N}$ is invertible in $R,$ with inverse in $z^{N}R\llbracket z\rrbracket.$ In particular, all integer powers $f^{n}\in R\llparenthesis z\rrparenthesis$ are defined in this case.
@@
Clearly, $f^{n}\cdot f^{m}=f^{n+m}$ by the binomial identity $\sum_{j=k+\ell}\binom{n}{k}\binom{m}{\ell}=\binom{n+m}{j}$ and $f^{n}\in z^{nN}R\llbracket z\rrbracket.$ Putting $n=1,m=-1,$ shows that $f(z)$ is invertible. ∎
@@ 3.2 Expansions
Expansion maps play a central role in modern formulations of vertex algebras. We generalize the classical theory and present a more conceptual approach here.
@@
Let $(x_{1},\ldots,x_{m}),\ldots,(z_{1},\ldots,z_{n})$ be groups of variables. The expansion maps are the inclusions of iterate Laurent rings
${R\llparenthesis x_{1},\ldots,x_{m}\rrparenthesis\cdots\llparenthesis z_{1},\ldots,z_{n}\rrparenthesis}$${R\llbracket x_{1}^{\pm 1},\ldots,x_{m}^{\pm 1},\cdots,z_{1}^{\pm 1},\ldots,z_{n}^{\pm 1}\rrbracket}$$\scriptstyle{i_{(x_{1},\ldots,x_{m}),\cdots,(z_{1},\ldots,z_{n})}}$
into the space of bilateral series. Expansion maps preserve products,
$i_{(x_{1},\ldots,x_{m}),\cdots,(z_{1},\ldots,z_{n})}(fg)=i_{(x_{1},\ldots,x_{m}),\cdots,(z_{1},\ldots,z_{n})}(f)\cdot i_{(x_{1},\ldots,x_{m}),\cdots,(z_{1},\ldots,z_{n})}(g).$
Singleton brackets will be dropped from the notation. We can also include holomorphic variables which we indicate by an underline.
@@
The grouping of variables is important, because the algebra in $R\llparenthesis z_{1}\rrparenthesis\llparenthesis z_{2}\rrparenthesis$ and $R\llparenthesis z_{2}\rrparenthesis\llparenthesis z_{1}\rrparenthesis$ is different. For example, the elements $i_{z_{1},z_{2}}(z_{1}\pm z_{2})^{n}$ and $i_{z_{2},z_{1}}(z_{1}\pm z_{2})^{n}$ are different in $R\llbracket z_{1}^{\pm 1},z_{2}^{\pm 1}\rrbracket$ when $n<0.$ The purpose of expansion maps is to specify the ring in which the algebraic operations are performed. For holomorphic variables the grouping is unimportant since
@@
Note that $R\llparenthesis z_{1},z_{2}\rrparenthesis=R\llparenthesis z_{1}\rrparenthesis\llparenthesis z_{2}\rrparenthesis\cap R\llparenthesis z_{2}\rrparenthesis\llparenthesis z_{1}\rrparenthesis$ is a proper intersection since, for example, $\sum_{n=0}^{\infty}z_{1}^{-n}z_{2}^{n}\in R\llparenthesis z_{1}\rrparenthesis\llparenthesis z_{2}\rrparenthesis\setminus R\llparenthesis z_{1},z_{2}\rrparenthesis.$ On the intersection we have
$i_{z_{1},z_{2}}f=i_{z_{2},z_{1}}f,\qquad\forall f\in R\llparenthesis z_{1},z_{2}\rrparenthesis.$ (3.5)
In particular, $R\llbracket z_{1}^{\pm 1},z_{2}^{\pm 1}\rrbracket$ is an $R\llparenthesis z_{1},z_{2}\rrparenthesis$-module and expansions are linear,
$i_{z_{1},z_{2}}(f\cdot g)=f\cdot i_{z_{1},z_{2}}(g),\qquad\forall f\in R\llparenthesis z_{1},z_{2}\rrparenthesis,\enskip g\in R\llbracket z_{1}^{\pm 1},z_{2}^{\pm 1}\rrbracket.$ (3.6)

### v2

@@
[…]
Vertex $F$-Algebras and Their Associated Lie Algebra
[…]
Vertex $\boldsymbol{F}$-Algebras and Their Associated Lie AlgebraThis paper is a contribution to the Special Issue on Recent Advances in Vertex Operator Algebras in honor of James Lepowsky. The full collection is available at https://www.emis.de/journals/SIGMA/Lepowsky.html
[…]
Markus UPMEIER
[…]
M. Upmeier
[…]
Department of Mathematics, University of Aberdeen, Fraser Noble Building,
[…]
Received April 18, 2025, in final form January 03, 2026; Published online January 15, 2026
[…]
Vertex $F$-algebras are a deformation of the concept of an ordinary vertex algebra in which the additive formal group law is replaced by an arbitrary formal group law $F$. The main theorem of this paper constructs a Lie algebra from a vertex $F$-algebra – for the additive formal group law, this extends Borcherds’ well-known construction for ordinary vertex algebras. Our construction involves the new concept of an $F$-residue and some other new algebraic concepts, which are deformations of familiar concepts for the special case of an additive formal group law.
[…]
vertex algebras; formal group laws; Lie algebras
[…]
17B69; 17B65
@@ 1 Introduction and results
A fundamental theorem of Borcherds [1], which has important applications to representation theory, states that every vertex algebra determines a Lie algebra. The main result of this paper extends this result and shows that every vertex $F$-algebra also determines a Lie algebra. However, the usual construction of the Lie bracket in the vertex algebra setting, via the residue of the state-to-field correspondence, does not directly apply to vertex $F$-algebras. To overcome this difficulty, we introduce a deformation of the concept of a residue, the $F$-residue $\operatorname{Res}_{z=0}^{F}$. Establishing the Lie algebra structure in the generalized setting requires significantly more technical work than in the classical case.
The following result is our main theorem, and it is proven in Section 4.3.
@@
Let $(V,\mathbbm{1},\mathcal{S},Y)$ be a vertex $F$-algebra (see Definition 4.1). Then
$a,b]=\operatorname{Res}_{z=0}^{F}Y(a,z)b{\rm d}z$
defines a Lie bracket on the quotient $V/\sum_{n\geqslant 1}\mathcal{S}^{(n)}(V)$.
Besides the potential applications to representation theory, there is further motivation for generalizing the concept of a vertex algebra stemming from enumerative geometry: recently, Joyce [8, 9] has shown that vertex algebras play a central role in enumerative geometry, where the Lie algebra associated to the ordinary homology $H_{*}(\mathcal{M})$ of a moduli stack $\mathcal{M}$ (for example, the stack of perfect complexes over a projective variety). The Lie bracket is used to formulate wall-crossing formulas of enumerative invariants. The study of enumerative invariants in generalized homology $E_{*}(\mathcal{M})$ is becoming increasingly popular (most notably, the theory of $K$-theoretic enumerative invariants). For complex oriented generalized homology theories $E_{*}$ with formal group law $F$, similar constructions show that $E_{*}(\mathcal{M})$ is naturally a vertex $F$-algebra. Indeed, in [4] the author and Gross proved that the generalized homology of an $H$-space (plus some extra data) naturally has a vertex $F$-algebra structure. This extends Joyce’s construction to generalized homology. While a comprehensive theory of wall-crossing has not yet been developed for generalized enumerative invariants (for progress in the case of $K$-theory, see Liu [13]), it may well be expected that the wall-crossing formula will use the Lie bracket constructed in this paper.
The proof of our main theorem, as well as a better understanding of vertex $F$-algebras, requires the development of various deformations of familiar concepts ($F$-binomial coefficients, $F$-delta distributions, $F$-residues, and $F$-hyperderivatives), developed in Section 3. These constructions, which appear to be new to the literature, provide the tools for generalizing results for ordinary vertex algebras to vertex $F$-algebras. Section 2 reviews some background of formal groups and establishes certain properties we will use later. In Section 4, we define vertex $F$-algebras (our definition is equivalent to that of Li [12]), prove some meromorphicity properties, and prove our main theorem.
We use the following notation.
  • $R$ a commutative ring with unit,
  • $R\big[z^{\pm 1}\big]$ the ring of Laurent polynomials,
  • $R\llbracket z\rrbracket$ the ring of formal power series $\sum_{i=0}^{\infty}a_{i}z^{i}$,
  • $R\big\llbracket z^{\pm 1}\big\rrbracket$ the abelian group of bilateral Laurent series $\sum_{i=-\infty}^{+\infty}a_{i}z^{i}$ (note that $R\big\llbracket z^{\pm 1}\big\rrbracket$ is not a ring, but only has a partially defined product),
  • R((z))R(\!(z)\!) the ring of meromorphic series having a pole at 0.
@@ 2 Background on formal groups
$F(z,w)=F(w,z)=z+w+O(zw),\qquad F(z,F(w,v))=F(F(z,w),v).$ (2.1)
There is a unique inverse $\iota(z)=-z+O\bigl(z^{2}\bigr)\in R\llbracket z\rrbracket$ such that $F(z,\iota z)=0$.
@@
Over $R=\mathbb{Z}[s]$, we have the formal group law
@@
We view $F_{s}$ as a family of group laws over the affine line. At $s=0$, it specializes to the additive group law $F_{a}$ and at $s=1$ to the multiplicative group law $F_{m}$. All polynomial formal group laws are of the form (2.2), and other examples must involve infinite series.
@@
Over the ring of modular forms $R=\mathbb{Z}\big[\frac{1}{2},\epsilon,\delta\big]$, we have the elliptic formal group law
$F_{ell}(z,w)=\frac{z\sqrt{S(w)}+w\sqrt{S(z)}}{1-\epsilon z^{2}w^{2}},\qquad S(z)=1-2\delta z^{2}+\epsilon z^{4}.$
For $\epsilon=\delta^{2}$ it specializes to the group law $\frac{z+w}{1+\delta zw}$ (for $\delta=1$ equivalent to $L$-genus) and for $\epsilon=0$ it specializes to
@@
Let $p$ be a prime number and fix a positive power $q=p^{h}$. Set $\phi_{p}(z)=z+\sum_{n\geqslant 1}p^{-n}z^{q^{n}}$. In [5], it is shown that
@@
$\theta_{F}=p_{F}(z){\rm d}z,\qquad p_{F}(z)\in R\llbracket z\rrbracket,$ (2.4)
satisfying $p_{F}(0)=1$, $F^{*}(\theta_{F})=\pi_{1}^{*}(\theta_{F})+\pi_{2}^{*}(\theta_{F})$. Equivalently,
$p_{F}(F(z,w))\bigl(F^{1,0}(z,w){\rm d}z+F^{0,1}(z,w){\rm d}w\bigr)=p_{F}(z){\rm d}z+p_{F}(w){\rm d}w.$ (2.5)
@@
$\iota^{*}(\theta_{F})=\theta_{F}.$ (2.6)
@@
$\phi(F(z,w))=\phi(z)+\phi(w),\qquad\phi(0)=0,\qquad\phi^{\prime}(0)=1.$ (2.7)
Indeed, taking $\partial/\partial w$ of (2.7) and setting $w=0$ we find that $\phi^{\prime}(z){\rm d}z=\theta_{F}$ is the invariant $1$-form which, since $\mathbb{Q}\subset R$, has a primitive $\theta_{F}=d\phi$. For this reason, we write $p_{F}(z)=\phi^{\prime}(z)$ with the caveat that $\phi(z)$ is only defined if $\mathbb{Q}\subset R$. The composition inverse $\phi^{-1}(x)\in R\llbracket x\rrbracket$ is called the exponential.
@@
For $s\neq 0$, the logarithm of $F_{s}$ is $s^{-1}\log(1+sz)$. For $F_{ell}$, the logarithm is the functional inverse of the indefinite elliptic integral $\int S(z)^{-1/2}{\rm d}z$.
@@
The logarithm in Example 2.4 is the series $\phi(z)$ defined there. For $p=q$, we have
$p_{F}(z)^{-1}=F^{0,1}(z,0)=\frac{1}{\phi^{\prime}(z)}=\prod_{k=1}^{\infty}\frac{1}{1+z^{p^{k}}}.$
Hence the coefficient of $z^{d}$ in $p_{F}(z)^{-1}$ is the sum of $(-1)^{k}$ over all partitions $p^{n_{1}}+\dots+p^{n_{k}}=d$ into positive powers of the prime $p$.
@@
$L\cong\mathbb{Z}[p_{1},p_{2},\dots]$ is a polynomial ring over the integers. In particular, $L$ is torsion-free.
@@
Substituting $w=z$, the first factor is $\psi(0)=1$ and the second is $\phi^{\prime}(z)$. ∎
3 Formal calculus and $\boldsymbol{F}$-residues
@@ 3.1 Bilateral and Laurent series
Let $R\big\llbracket z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}\big\rrbracket$ be the space of bilateral Laurent series
$f=\sum_{i_{1},\dots,i_{n}\in\mathbb{Z}}a_{i_{1},\dots,i_{n}}z_{1}^{i_{1}}\cdots z_{n}^{i_{n}}$ (3.1)
in variables $z_{1},\dots,z_{n}$. This is an abelian group under addition, but the product is only partially defined. The product of (3.1) with
$g=\sum_{j_{1},\dots,j_{n}\in\mathbb{Z}}b_{j_{1},\dots,j_{n}}z_{1}^{j_{1}}\cdots z_{n}^{j_{n}}$
@@
$fg=\sum_{k_{1},\dots,k_{n}\in\mathbb{Z}}\Biggl(\sum_{i_{1}+j_{1}=k_{1},\dots,i_{n}+j_{n}=k_{n}}a_{i_{1},\dots,i_{n}}b_{j_{1},\dots,j_{n}}\Biggr)z_{1}^{k_{1}}\cdots z_{n}^{k_{n}}$
reduces to a sum with only finitely many non-zero terms. By [11, p. 24], this product is associative if all products $fg$, ghgh, $(fg)h$, $f(gh)$ converge and the triple product converges.
@@
We say that (3.1) is a formal Laurent series if $a_{i_{1},\dots,i_{n}}=0$ for all but finitely many negative indices. The subspace of formal Laurent series $R(\!(z_{1},\dots,z_{n})\!)$ is a ring since all products converge. For $f\in R(\!(z_{1},\dots,z_{n})\!)$, we say also that $f$ is meromorphic in the variables $z_{1},\dots,z_{n}$. We have the subring $R\llbracket\underline{z}_{1},\dots,\underline{z}_{n}\rrbracket$ of formal power series which are also said to be holomorphic in $\underline{z}_{1},\dots,\underline{z}_{n}$. It is useful to underline holomorphic variables in some contexts below.
@@
This generalizes to series with coefficients in an $R$-module $M$. Then $R(\!(z_{1},\dots,z_{n})\!)$ acts on $M(\!(z_{1},\dots,z_{n})\!)$. We leave this extension to the reader.
@@
Let $f=\sum_{n\geqslant N}a_{n}z^{n}\in R(\!(z)\!)$ with lowest coefficient $a_{N}\neq 0$. Then $f$ is invertible in R((z))R(\!(z)\!) if and only if $a_{N}$ is invertible in $R$, with inverse in $z^{N}R\llbracket z\rrbracket$. In particular, all integer powers $f^{n}\in R(\!(z)\!)$ are defined in this case.
@@
Clearly, $f^{n}\cdot f^{m}=f^{n+m}$ by the binomial identity $\sum_{j=k+\ell}\binom{n}{k}\binom{m}{\ell}=\binom{n+m}{j}$ and $f^{n}\in z^{nN}R\llbracket z\rrbracket$. Putting $n=1$, m=1m=-1, shows that $f(z)$ is invertible. ∎
@@ 3.2 Expansions
Expansion maps play a central role in modern formulations of vertex algebras, so we briefly review them here.
@@
Let $(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})$ be tuples of formal variables, where singleton brackets will be dropped from the notation. We view the iterate Laurent series ring as a subset of the space of bilateral series:
$R(\!(x_{1},\dots,x_{m})\!)\cdots(\!(z_{1},\dots,z_{n})\!)\subset R\big\llbracket x_{1}^{\pm 1},\dots,x_{m}^{\pm 1},\dots,z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}\big\rrbracket.$
Let $j_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}$ be the natural inclusion of $R(\!(x_{1},\dots,x_{m},\dots,z_{1},\dots,z_{n})\!)$, the un-iterated Laurent ring, into the ring $R(\!(x_{1},\dots,x_{m})\!)\cdots(\!(z_{1},\dots,z_{n})\!)$. Moreover, let $S_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}\subset R(\!(x_{1},\dots,x_{m},\dots,z_{1},\dots,z_{n})\!)$ be the multiplicative set consisting of those Laurent series whose image under $j_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}$ is invertible in the iterated Laurent ring.
The expansion map $i_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}$ is the localization of the map $j_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}$ at $S_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}$} shown in the diagram
$\thinspace R(\!(x_{1},\dots,x_{m},\dots,z_{1},\dots,z_{n})\!)$$R(\!(x_{1},\dots,x_{m})\!)\cdots(\!(z_{1},\dots,z_{n})\!)$$S_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}^{-1}R(\!(x_{1},\dots,x_{m},\dots,z_{1},\dots,z_{n})\!).$$\scriptstyle{j_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}}$$\scriptstyle{i_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}}$
Example 3.4.
Let $F(z,w)$ be a formal group law. The images of the series $F(z,w)$, $F(z,\iota w)$, $F(\iota z,w)$, $\iota F(z,w)$ in R((z))((w))R(\!(z)\!)(\!(w)\!) are invertible because their lowest coefficients are the units $z,\iota(z)$ in R((z))R(\!(z)\!). Therefore, $F(z,w)$, $F(z,\iota w)$, $F(\iota z,w)$, $\iota F(z,w)$ are in $S_{z,w}$. Hence there are well-defined integer powers
$i_{z,w}F(z,w)^{n},\qquad i_{z,w}F(\iota z,w)^{n},\qquad i_{z,w}F(z,\iota w)^{n},\qquad i_{z,w}(\iota F(z,w))^{n},\qquad\forall n\in\mathbb{Z},$
which are elements of $R(\!(z)\!)(\!(w)\!)\subset R\big\llbracket z^{\pm 1},w^{\pm 1}\big\rrbracket$. These are computed by first viewing $F(z,w)$, $F(z,\iota w)$, $F(\iota z,w)$, $\iota F(z,w)$ as elements of the iterated ring R((z))((w))R(\!(z)\!)(\!(w)\!) and then forming the $n$-th power there.
@@
For ordinary vertex algebras, the expansion maps $R(\!(z,w)\!)[z-w]^{-1}\to R(\!(z)\!)(\!(w)\!)$ are defined on the localization by a single element zwz-w. For vertex $F$-algebras, it becomes necessary to localize $F(z,w)$, $F(z,\iota w)$, $F(\iota z,w)$, $\iota F(z,w)$. In the case of several variables, even more complicated expressions in $F$ must be localized (for example, in the proof of Proposition 3.19). Since these expressions are difficult to list systematically, we define the expansion maps on the universal localization, for example, $i_{z,w}\colon S_{z,w}^{-1}R(\!(z,w)\!)\to R(\!(z)\!)(\!(w)\!)$.
We can also include holomorphic variables which we indicate by an underline. Note that expansion maps preserve products,
$i_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}(fg)=i_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}(f)\cdot i_{(x_{1},\dots,x_{m}),\dots,(z_{1},\dots,z_{n})}(g).$
Remark 3.6.
The grouping of variables is important, because the algebra in $R(\!(z_{1})\!)(\!(z_{2})\!)$ and $R(\!(z_{2})\!)(\!(z_{1})\!)$} is different. For example, the elements $i_{z_{1},z_{2}}(z_{1}\pm z_{2})^{n}$ and $i_{z_{2},z_{1}}(z_{1}\pm z_{2})^{n}$ are different in $R\big\llbracket z_{1}^{\pm 1},z_{2}^{\pm 1}\big\rrbracket$ when $n<0$. Expansion maps specify the ring in which the algebraic operations are performed. For holomorphic variables, the grouping is unimportant since
@@
Note that $R(\!(z_{1},z_{2})\!)=R(\!(z_{1})\!)(\!(z_{2})\!)\cap R(\!(z_{2})\!)(\!(z_{1})\!)$ is a proper intersection since, for example, $\sum_{n=0}^{\infty}z_{1}^{-n}z_{2}^{n}\in R(\!(z_{1})\!)(\!(z_{2})\!)\setminus R(\!(z_{1},z_{2})\!)$. On the intersection, we have
$i_{z_{1},z_{2}}f=i_{z_{2},z_{1}}f,\qquad\forall f\in R(\!(z_{1},z_{2})\!).$ (3.3)
In particular, $R\big\llbracket z_{1}^{\pm 1},z_{2}^{\pm 1}\big\rrbracket$ is an $R(\!(z_{1},z_{2})\!)$-module and expansions are linear,
$i_{z_{1},z_{2}}(f\cdot g)=f\cdot i_{z_{1},z_{2}}(g),\qquad\forall f\in R(\!(z_{1},z_{2})\!),\ g\in R\big\llbracket z_{1}^{\pm 1},z_{2}^{\pm 1}\big\rrbracket.$ (3.4)